# POINTS: Improving Your Vision-language Model with Affordable Strategies

### v2

@@ 3. $\bullet$
$f(\theta_{s})$ $=f(\frac{1}{p}\sum_{i=1}^{k}\theta_{i})$ (5)
@@ 4 Experiments
This section is divided into five subsections: (i) evaluation setup, (ii) pre-training and instruction-tuning datasets used to train the strong baseline, as well as the instruction-tuning datasets selected by model soup, (iii) details about the training setup for the OCR ViT pre-training, the vision-language pre-training, and the visual instruction tuning stages, (iv) ablation studies and analyses of each component used to build our final model, and (v) comparison with other works on extensive benchmarks.
@@ 4.1 Evaluation Setup
Before embarking on our exploration, we sought a robust evaluation metric to comprehensively assess the various capabilities of our model. This is where OpenCompass (Contributors, 2023) proves helpful. OpenCompass proposes eight benchmarks to balance the evaluation of a model from different perspectives. These benchmarks include MMBench (Liu et al., 2023c) and MMStar (Chen et al., 2024a) for diagnosing general abilities, MMMU (Yue et al., 2024) for testing STEM-related abilities, HallusionBench (Liu et al., 2023a) for model hallucination, MathVista (Lu et al., 2023) for math-related abilities, AI2D (Kembhavi et al., 2016) for chart-related abilities, OCRBench (Liu et al., 2023d) for OCR capabilities, and MMVet (Yu et al., 2023) for subjective evaluation. By averaging the metrics from these benchmarks, OpenCompass derives a score that represents the comprehensive ability of a model. Additionally, it offers a useful tool, VLMEvalKit (Duan et al., 2024), for one-click evaluation. Therefore, unless otherwise specified, we will use these eight benchmarks for our ablation study, with the exception of MMBench, for which we will use the dev-en split.
@@ 4.2 Data Setup
To train the OCR ViT, we randomly selected 20 million data points from LAION-5B-en (Schuhmann et al., 2022), LAION-5B-cn (Schuhmann et al., 2022), WuKong (Gu et al., 2022), and Zero (Gu et al., 2022). We then used PaddleOCR to extract text from the images, replacing the original captions to form new image-caption pairs for pre-training. Following Vary (Wei et al., 2023), we also included 10 million original data samples from LAION-5B, where the captions are the original ones crawled from the Internet. However, we did not adopt the cumbersome pipeline of constructing a new dataset for OCR enhancement, such as crawling PDF files and converting them to images for training (Bai et al., 2023b), as we found our existing pipeline already performs well on OCR-related tasks. For the vision-language pre-training in constructing the strong baseline, we used CapFusion to construct 20 million data points (note that these data do not overlap with those used in OCR ViT pre-training) from LAION-5B. From this set, we selected 5 million data points, as we found this setting works best, similar to the observation in Liu et al. (2024c). Based on the 5 million data points, we further selected a 1 million dataset for the final vision-language alignment by choosing the top 20% of data with the lowest perplexity value.
@@
Based on the datasets identified by Liu et al. (2024c), we further employ Individual Select to choose additional datasets from those proposed in (Lu et al., 2024a), (Tong et al., 2024), and (Laurençon et al., 2024). The final composition of datasets, referred to as the Base Set, used to construct the robust baseline is presented in appendix. Additionally, we include the datasets selected by model soup to derive the final model, which are highlighted in red.
@@ 4.3 Training Setup
The pre-training framework follows the standard LLaVA-style architecture (Liu et al., 2023b), comprising a vision encoder, a two-layer MLP, and a large language model. The vision encoder is initialized from OpenAI’s CLIP-ViT-Large-336111https://huggingface.co/openai/clip-vit-large-patch14-336, while the large language model is initialized from Yi-1.5-9B-Chat (Young et al., 2024). Throughout the pre-training stage, the large language model remains frozen, whereas the vision encoder and MLP are trainable. The learning rates for the vision encoder and MLP are set to $2\times 10^{-4}$ and $2\times 10^{-5}$, respectively, with a warm-up schedule during the first 3% of steps, followed by a cosine decay schedule for the remaining steps.
@@ 4.4 Ablation Study and Analysis
As shown in Table 1, each component introduced in subsection 3.1 contributes to steady improvements. These enhancements are significant; for instance, after introducing Dynamic High Resolution to split the input image, we observe substantial improvements in OCR-related tasks, such as OCRBench, with performance increasing from 49.6% to 55.6%. Additionally, the use of high-resolution images with Dynamic High Resolution helps reduce hallucination, primarily due to the increased detail in the high-resolution images. Furthermore, replacing the original Dynamic High Resolution with CATTY results in notable improvements across various benchmarks, with OCR-related benchmarks showing greater gains than others. This is likely because image distortion has a more pronounced negative impact on text within images. Compared to general visual feature extraction, the ability to extract text features from images is limited for CLIP-ViT(Radford et al., 2021), as it was trained on a large quantity of general image-text pairs. Consequently, we observe substantial improvements on OCRBench after integrating features from an additional ViT, post-trained on text-rich images. Among the 5 strategies, incorporating more visual instruction tuning datasets by Individual Select yields the most significant improvements. This observation aligns with existing works(Chen et al., 2024c; Li et al., 2024; Tong et al., 2024), underscoring the importance of selecting effective datasets during the visual instruction tuning stage.
@@
Table 1: Ablation about each component to build the strong baseline. CF: CapFusion(Yu et al., 2024), DHR: Dynamic high resolution(Chen et al., 2024c), CATTY: Consistent aspect ratio dynamic high resolution proposed by us, DEV: Dual vision encoder(Wei et al., 2023), IS: Individual select(Liu et al., 2024c). MMB: the dev-en split of MMBench(Liu et al., 2023c), MV: MathVista(Lu et al., 2023), HB: HallusionBench(Liu et al., 2023a), OCR: OCRBench(Liu et al., 2023d), Overall: the average of scores on the first 8 benchmarks.
@@
Table 3: DVE: Dual Vision Encoder architecture. SVE: Single Vision Encoder architecture, incorporating the OCR dataset used to train the OCR ViT into the dataset for the vision-language pre-training stage.
@@
Pre-train Dataset
As shown in Table 3, scaling up the dataset size (constructed by CapFusion) from 5M to 20M results in downgraded performance, similar to the observations in Liu et al. (2024c). Additionally, some works also achieve promising performance using relative small pre-training datasets instead of a huge number of datasets during the pre-training stage (Li et al., 2024; Liu et al., 2024a). We believe the possible reasons are: i) The vision encoder of most existing vision-language models is initialized from a pre-trained model that has already been trained on a large quantity of image-text pairs. It is highly likely that most of the data used in the vision-language pre-training stage has already been seen by the vision encoder, thus bringing only marginal or even negative impact when scaling up the size of the vision-language pre-training dataset. ii) The pre-training datasets are quite homogeneous for existing large-scale web-crawled datasets, e.g., LAION-5B and COYO-700M (Byeon et al., 2022). We plot the distribution of the main entity for each image of a subset extracted from LAION-5B in the appendix and find that this distribution is long-tailed and constrained to a few objects, e.g., person. Thus, indiscriminately pre-training the model on such datasets can only bring limited benefits. As shown in the third row, we can improve performance by pre-training the model on merely 1M data, coming from the top 20% of the 5M data in the first row, which has the lowest perplexity. This result shows that excessively exposing the model to obscure and scarce knowledge during the transition is detrimental to its learning. Furthermore, compared to fusing features from a separate OCR-enhanced vision encoder, introducing a large OCR dataset has two obvious drawbacks: i) During the pre-training stage, the model has to align both the general features and OCR-related features, which may result in conflicts (Wei et al., 2023). ii) Since the size of the dataset used in vision-language pre-training is relatively small, a large OCR dataset may overwhelm the learning process, which is not helpful for learning other kinds of knowledge. Thus, introducing features from another OCR ViT can yield superior performance in Table 4.
@@
Table 5: Comparison between different methods. MMB: the dev-en split of MMBench(Liu et al., 2023c), MV: MathVista(Lu et al., 2023), HB: HallusionBench(Liu et al., 2023a), OCR: OCRBench(Liu et al., 2023d), SCI: ScienceQA(Lu et al., 2022), MME: MME(Yin et al., 2023), RWQ: RealWorldQA, Wild: LLaVA-Wild(Liu et al., 2024b). Cambrian-34: Cambrian-34B(Tong et al., 2024), Ovis1.5-LLaMA3-8B: Ovis1.5(Lu et al., 2024b), IXC-2.5: InternLM-XComposer-2.5(Zhang et al., 2024), OneVision: LLaVA-OneVision(Li et al., 2024). The language model POINTS-9B uses is Yi-1.5-9B (Young et al., 2024).
@@
As described in previous sections, increasing adding more instruction tuning datasets often reaches a plateau where further increasing the number of datasets yields minimal improvement. However, by incorporating model soup across different datasets, we observe substantial enhancements, as shown in Table 5, with the overall score increasing from 59.0 to 61.8. We also compare the benefits of various model soup strategies. Among them, greedy soup achieves the best performance, outperforming maximum soup and average soup by 0.8 and 0.6 points, respectively. Unless otherwise specified, we will use greedy soup by default in subsequent experiments. Additionally, we include the results of conducting model soup over different hyperparameters, e.g. different learning rates. As shown, model soup over hyperparameters brings only marginal improvement; detailed results are provided in the appendix. Furthermore, we verify in the appendix that model soup consistently improves performance regardless of the Base Set used.
@@ 4.5 Comparison with Other Works
In addition to the 8 benchmarks used in the ablation studies above, we further include ScienceQA (Lu et al., 2022), MME (Yin et al., 2023), LLaVA-Wild (Liu et al., 2024b), and ReadWorldQA to compare the performance of different models. The following table shows the performance of these models. As shown in Table 5, POINTS achieves performance comparable to existing state-of-the-art models of similar size and even surpasses models with much larger sizes, such as Cambrian-34B. Additionally, compared to the models listed in the table, POINTS uses a much smaller pre-training dataset (e.g., 1M), fewer visual instruction tuning datasets, and all the datasets we used are publicly available. This makes it more affordable for the community to adopt the strategies proposed in this paper. Furthermore, each aspect of POINTS is clearly presented and thoroughly analyzed, making the effectiveness of each strategy employed in our model evident.
@@ 5 Conclusion
Vision-language models have achieved significant progress in recent years. Following this trend (Chen et al., 2024c; Li et al., 2024; Liu et al., 2024b; Zhang et al., 2024; Tong et al., 2024), we first establish a strong baseline by integrating various advancements proposed in recent works (Liu et al., 2024a; Yu et al., 2024; Wei et al., 2023; Liu et al., 2024c) for further experiments. Additionally, we delve into the intricate details of these advancements and propose effective refinements, such as the Consistent Aspect Ratio Dynamic High Resolution. We also conduct extensive experiments to verify the effectiveness of each component in constructing the strong baseline. Secondly, we propose using perplexity to filter the pre-training dataset, retaining only the top 20% of data with the smallest perplexity values during the pre-training stage. This filtering method also brings significant improvements. Model Soup (Wortsman et al., 2022) has shown promising potential to further enhance performance by averaging the weights of fine-tuned models with different hyperparameters. However, we find that conducting model soup over different dataset settings can yield even more substantial improvements. In this paper, we propose conducting model soup across models fine-tuned with varying dataset quantities and diversities. The improvements brought by quantity and diversity are orthogonal to each other and can result in joint enhancement.

### v3

@@ 3. $\bullet$
$f(\theta_{s})$ $=f(\frac{1}{k}\sum_{i=1}^{k}\theta_{i})$ (5)
@@ 4 Experiments
This section is divided into five subsections: (i) evaluation setup, (ii) pre-training and instruction-tuning datasets used to train the strong baseline (iii) details about the training setup for the OCR ViT pre-training, the vision-language pre-training, and the visual instruction tuning stages, (iv) ablation studies and analyses of each component used to build our final model, and (v) comparison with other works on extensive benchmarks.
@@ 4.1 Evaluation Setup
Before embarking on our exploration, we sought a robust evaluation metric to comprehensively assess the various capabilities of our model. This is where OpenCompass (Contributors, 2023) proves helpful. OpenCompass proposes eight benchmarks to balance the evaluation of a model from different perspectives. These benchmarks include MMBench (Liu et al., 2023c) and MMStar (Chen et al., 2024b) for diagnosing general abilities, MMMU (Yue et al., 2024) for testing STEM-related abilities, HallusionBench (Liu et al., 2023a) for model hallucination, MathVista (Lu et al., 2023) for math-related abilities, AI2D (Kembhavi et al., 2016) for chart-related abilities, OCRBench (Liu et al., 2023d) for OCR capabilities, and MMVet (Yu et al., 2023b) for subjective evaluation. By averaging the metrics from these benchmarks, OpenCompass derives a score that represents the comprehensive ability of a model. Additionally, it offers a useful tool, VLMEvalKit (Duan et al., 2024), for one-click evaluation. Therefore, unless otherwise specified, we will use these eight benchmarks for our ablation study, with the exception of MMBench, for which we will use the dev-en split.
@@ 4.2 Data Setup
To train the OCR ViT, we randomly selected 20 million data points from LAION-5B-en (Schuhmann et al., 2022), LAION-5B-cn (Schuhmann et al., 2022), WuKong (Gu et al., 2022), and Zero (Gu et al., 2022). We then used PaddleOCR to extract text from the images, replacing the original captions to form new image-caption pairs for pre-training. Following Vary (Wei et al., 2023), we also included 10 million original data samples from LAION-5B, where the captions are the original ones crawled from the Internet. However, we did not adopt the cumbersome pipeline of constructing a new dataset for OCR enhancement, such as crawling PDF files and converting them to images for training (Bai et al., 2023b), as we found our existing pipeline already performs well on OCR-related tasks. For the vision-language pre-training in constructing the strong baseline, we used CapFusion to construct 20 million data points (note that these data do not overlap with those used in OCR ViT pre-training) from LAION-5B. From this set, we selected 5 million data points, as we found this setting works best, similar to the observation in Liu et al. (2024c). Based on the 5 million data, we further selected a 1 million dataset for the final vision-language alignment by choosing the top 20% of data with the lowest perplexity value.
@@
Based on the datasets identified by Liu et al. (2024c), we further employ Individual Select to choose additional datasets from those proposed in (Lu et al., 2024a), (Tong et al., 2024), and (Laurençon et al., 2024b). The final composition of datasets, referred to as the Base Set, used to construct the robust baseline is presented in Appendix.
@@ 4.3 Training Setup
The pre-training framework follows the standard LLaVA-style architecture (Liu et al., 2023b), comprising a vision encoder, a two-layer MLP, and a large language model. The vision encoder is initialized from OpenAI’s CLIP-ViT-Large-336, while the large language model is initialized from Yi-1.5-9B-Chat (Young et al., 2024). Throughout the pre-training stage, the large language model remains frozen, whereas the vision encoder and MLP are trainable. The learning rates for the vision encoder and MLP are set to $2\times 10^{-4}$ and $2\times 10^{-5}$, respectively, with a warm-up schedule during the first 3% of steps, followed by a cosine decay schedule for the remaining steps.
@@ 4.4 Ablation Study and Analysis
As shown in Table 1, each component introduced in subsection 3.1 contributes to steady improvements. These enhancements are significant; for instance, after introducing Dynamic High Resolution to split the input image, we observe substantial improvements in OCR-related tasks, such as OCRBench, with performance increasing from 49.6% to 55.6%. Additionally, the use of high-resolution images with Dynamic High Resolution helps reduce hallucination, primarily due to the increased detail in the high-resolution images. Furthermore, replacing the original Dynamic High Resolution with CATTY results in notable improvements across various benchmarks, with OCR-related benchmarks showing greater gains than others. This is likely because image distortion has a more pronounced negative impact on text within images. Compared to general visual feature extraction, the ability to extract text features from images is limited for CLIP-ViT(Radford et al., 2021), as it was trained on a large quantity of general image-text pairs. Consequently, we observe substantial improvements on OCRBench after integrating features from an additional ViT, post-trained on text-rich images. Among the 5 strategies, incorporating more visual instruction tuning datasets by Individual Select yields the most significant improvements. This observation aligns with existing works(Chen et al., 2024d; Li et al., 2024a; Tong et al., 2024), underscoring the importance of selecting effective datasets during the visual instruction tuning stage.
@@
Table 1: Ablation about each component to build the strong baseline. CF: CapFusion(Yu et al., 2024), DHR: Dynamic high resolution(Chen et al., 2024d), CATTY: Consistent aspect ratio dynamic high resolution proposed by us, DEV: Dual vision encoder(Wei et al., 2023), IS: Individual select(Liu et al., 2024c). MMB: the dev-en split of MMBench(Liu et al., 2023c), MV: MathVista(Lu et al., 2023), HB: HallusionBench(Liu et al., 2023a), OCR: OCRBench(Liu et al., 2023d), Overall: the average of scores on the first 8 benchmarks.
@@
Table 3: DVE: Dual Vision Encode. SVE: Single Vision Encoder, incorporating the OCR dataset used to train the OCR ViT into the dataset for the vision-language pre-training stage.
@@
POINTS-7B 83.2 63.1 46.0 72.0 80.9 52.3 61.0 49.4 94.8 2195.2 67.3 71.1
Table 5: Comparison between different methods. MMB: the dev-en split of MMBench(Liu et al., 2023c), MV: MathVista(Lu et al., 2023), HB: HallusionBench(Liu et al., 2023a), OCR: OCRBench(Liu et al., 2023d), SCI: ScienceQA(Lu et al., 2022a), MME: MME(Yin et al., 2023), RWQ: RealWorldQA, Wild: LLaVA-Wild(Liu et al., 2024b). Cambrian-34: Cambrian-34B(Tong et al., 2024), Ovis1.5-LLaMA3-8B: Ovis1.5(Lu et al., 2024b), IXC-2.5: InternLM-XComposer-2.5(Zhang et al., 2024), OneVision: LLaVA-OneVision(Li et al., 2024a), Idefics3-LLaMA3-8B: IDEFICS3 (Laurençon et al., 2024a). The language model POINTS-9B uses is Yi-1.5-9B (Young et al., 2024). Results are obtained from the leaderboard of OpenCompass, except for MMBench. POINTS-7B uses is Qwen-2.5-7B (Team, 2024).
Pre-train Dataset
As shown in Table 2, scaling up the dataset size (constructed by CapFusion) from 5M to 20M results in downgraded performance, similar to the observations in Liu et al. (2024c). Additionally, some works also achieve promising performance using relative small pre-training datasets instead of a huge number of datasets during the pre-training stage (Li et al., 2024a; Liu et al., 2024a). We believe the possible reasons are: i) The vision encoder of most existing vision-language models is initialized from a pre-trained model that has already been trained on a large quantity of image-text pairs. It is highly likely that most of the data used in the vision-language pre-training stage has already been seen by the vision encoder, thus bringing only marginal or even negative impact when scaling up the size of the vision-language pre-training dataset. ii) The pre-training datasets are quite homogeneous for existing large-scale web-crawled datasets, e.g., LAION-5B and COYO-700M (Byeon et al., 2022). We plot the distribution of the main entity for each image of a subset extracted from LAION-5B in the Appendix and find that this distribution is long-tailed and constrained to a few objects, e.g., person. Thus, indiscriminately pre-training the model on such datasets can only bring limited benefits. As shown in the third row, we can improve performance by pre-training the model on merely 1M data, coming from the top 20% of the 5M data in the first row, which has the lowest perplexity. This result shows that excessively exposing the model to obscure and scarce knowledge during the transition is detrimental to its learning. Furthermore, compared to fusing features from a separate OCR-enhanced vision encoder, introducing a large OCR dataset has two obvious drawbacks: i) During the pre-training stage, the model has to align both the general features and OCR-related features, which may result in conflicts (Wei et al., 2023). ii) Since the size of the dataset used in vision-language pre-training is relatively small, a large OCR dataset may overwhelm the learning process, which is not helpful for learning other kinds of knowledge. Thus, introducing features from another OCR ViT can yield superior performance in Table 4.
@@
As described in previous sections, increasingly adding more instruction tuning datasets often reaches a plateau where further increasing the number of datasets yields minimal improvement. However, by incorporating model soup across different datasets, we observe substantial enhancements, as shown in Figure 3, with the overall score increasing from 59.0 to 61.8. We also compare the benefits of various model soup strategies. Among them, greedy soup achieves the best performance, outperforming maximum soup and average soup by 0.8 and 0.6 points, respectively. Unless otherwise specified, we will use greedy soup by default in subsequent experiments. Additionally, we include the results of conducting model soup over different hyperparameters, e.g. different learning rates. As shown, model soup over hyperparameters brings only marginal improvement. Furthermore, we verify in the Appendix that model soup consistently improves performance regardless of the Base Set used.
@@ 4.5 Comparison with Other Works
In addition to the 8 benchmarks used in the ablation studies above, we further include ScienceQA (Lu et al., 2022a), MME (Yin et al., 2023), LLaVA-Wild (Liu et al., 2024b), and ReadWorldQA to compare the performance of different models. The following table shows the performance of these models. As shown in Table 5, POINTS achieves performance comparable to existing state-of-the-art models of similar size and even surpasses models with much larger sizes, such as Cambrian-34B. Additionally, compared to the models listed in the table, POINTS uses a much smaller pre-training dataset (e.g., 1M), fewer visual instruction tuning datasets, and all the datasets we used are publicly available. This makes it more affordable for the community to adopt the strategies proposed in this paper. Furthermore, each aspect of POINTS is clearly presented and thoroughly analyzed, making the effectiveness of each strategy employed in our model evident.
@@ 5 Conclusion
Vision-language models have achieved significant progress in recent years. Following this trend (Chen et al., 2024d; Li et al., 2024a; Liu et al., 2024b; Zhang et al., 2024; Tong et al., 2024), we first establish a strong baseline by integrating various advancements proposed in recent works (Liu et al., 2024a; Yu et al., 2024; Wei et al., 2023; Liu et al., 2024c) for further experiments. Additionally, we delve into the intricate details of these advancements and propose effective refinements, such as the Consistent Aspect Ratio Dynamic High Resolution. We also conduct extensive experiments to verify the effectiveness of each component in constructing the strong baseline. Secondly, we propose using perplexity to filter the pre-training dataset, retaining only the top 20% of data with the smallest perplexity values during the pre-training stage. This filtering method also brings significant improvements. Model Soup (Wortsman et al., 2022) has shown promising potential to further enhance performance by averaging the weights of fine-tuned models with different hyperparameters. However, we find that conducting model soup over different dataset settings can yield even more substantial improvements.
Appendix A Visual Instruction Tuning Datasets
Table 6 shows the visual instruction tuning datasets used in this work.
Category Dataset
Conversation LLaVAR(Zhang et al., 2023), inhouse GPT-4o data, Mini-Gemini(Li et al., 2024b)
LVIS-Instruct4V(Wang et al., 2023)
Document DocVQA(en)(Mathew et al., 2021)
Caption ALLaVA(Chen et al., 2024a), ShareGPT4V(Chen et al., 2023), LAION-GPT4V
General QA VSR(Zhang et al., 2021), IConQA(Lu et al., 2021b)
Science AI2D(Kembhavi et al., 2016), TQA(Kim et al., 2018), ScienceQA(Lu et al., 2022a)
Chart&Screen DVQA(Kafle et al., 2018), POIE(Kuang et al., 2023), MapQA(Chang et al., 2022)
ScreenQA(Hsiao et al., 2022)
Mathematics GeoQA+(Cao & Xiao, 2022), Geo3K(Lu et al., 2021a), TabMWP(Lu et al., 2022b)
CLEVR-Math(Lindström & Abraham, 2022), SuperCLEVER(Li et al., 2023b),
MathV360K(Shi et al., 2024)
Knowledge KVQA(Sanket Shah & Talukdar, 2019)
OCR InfoVQA(Mathew et al., 2022), TextVQA(Singh et al., 2019)
ST-VQA(Biten et al., 2019), ICDAR2015, HME100K(Yuan et al., 2022)
Text-only LIMA(Zhou et al., 2024), Alpaca-GPT4(Peng et al., 2023)
OpenHermes2.5(Teknium, 2023), MetaMathQA(Yu et al., 2023a)
MathInstruct(Yue et al., 2023), orca-math-word-problems-200k(Mitra et al., 2024)
atlas-math-sets, Math
Table 6: Visual instruction tuning datasets to build the strong baseline and the those finally selected to conduct model soup (marked in red).
Appendix B Prompt for CapFusion
Figure 4: Prompt for image caption generation and captions merging.
Figure 4 shows the prompts to generate new image caption and merge the original caption and the generated caption.
Appendix C Prompt for inhouse GPT-4o Dataset
Figure 5 shows the prompt to generate the inhouse GPT-4o data in Table 6.
Figure 5: Prompt to generate the inhouse GPT-4o dataset.
Appendix D Model Soup with Different Base Set
To verify that model soup consistently improves performance regardless of the Base Set used, we randomly sampled 6 datasets from the Base Set in Table 6 to conduct model soup, while the remaining datasets were used as the new Base Set. As shown in Figure 6, model soup also brings significant improvements compared to individual selection, demonstrating the effectiveness and universality of model soup.
Figure 6: Model Soup brings consistent improvement regardless of what Base Set is used.
Appendix E Homogeneity in Existing Pre-training Dataset
We randomly sampled 5 million images from LAION-5B and used POINTS to identify the main object in each image. We then plotted the distribution of the top six objects from the sampled data. As shown on the left side of Figure 7, these six objects account for more than 90% of the total data. The right side of Figure 7 illustrates the distribution of the top six objects after applying a simple balancing technique: if the count of a particular object exceeds the average count of the top six objects, we down-sample it to 60% of its total count. We re-trained the strong baseline model on both the original and balanced pre-training datasets. The overall score of the balanced version outperformed the original by 0.6. This is an initial investigation into the distribution of the pre-training dataset, and we plan to explore this direction further in our future research.
Figure 7: Top 6 objects from the subset we randomly sample from LAION-5B (left), and the distribution of the top 6 objects after simple balance (right).
Appendix F Case Study of Pre-training Dataset
As discussed in Section 3.2, we filter the pre-training dataset using perplexity. Figure 8 shows samples randomly selected from the top 20% and bottom 20% of the data, which have the lowest perplexity values. As illustrated, the captions in the bottom 20% of the data are more likely to contain obscure world knowledge. While augmenting the model with more world knowledge during pre-training can help it generalize better in real-world scenarios, the relatively small scale of the vision-language model pre-training dataset makes this obscure world knowledge in the bottom 20% quite sparse (seldom appearing more than once during pre-training). Consequently, this world knowledge is more likely to be noise rather than informative content for pre-training. Additionally, we also use InternVL2 (Chen et al., 2024d) to filter the pre-training dataset. The model pre-trained on the filtered subset achieves an overall score of 60.1, surpassing the model pre-trained on the original 5M dataset by 1.1 points.
Figure 8: The left are samples randomly selected from the top 20% and the right are samples randomly selected from the last 20%. These obscure world knowledge is marked in red.
Appendix G Ablation about the Maximum Number of Tiles
We perform more fine-grained ablation studies about the maximum number of tiles used in CATTY, and Figure 9 shows the results.
Figure 9: Ablation study about the maximum split used in CATTY.
Appendix H Ablation about the Feature Average from Vision Encoder
Before feeding features into the LLM, we compute the weighted average of features from both the general and OCR vision encoders. Figure 10 illustrates the model’s performance when different weights are assigned to the general vision encoder (note that the weights assigned to the general and OCR vision encoders sum to 1). In this ablation study, we adhere to the experimental settings described in the fifth row of Table 1 in the main paper.
Figure 10: Ablation about the weight assigned to the general vision encoder.
Appendix I Training Cost of POINTS
[…]